# Biased Minds Meet Biased AI: How Class Imbalance Shapes Appropriate Reliance and Interacts with Human Base Rate Neglect

### v1

@@ 7.2.4. Broader Implications for AI Research and Practice
Integrating BRN within the broader framework of human biases, our findings may support the view that BRN involves deliberate processing (system two) rather than being purely automatic and intuitive (pennycook_base_2014; MARKOVITS2023105451). A plausible cognitive mechanism is that AI advice introduces additional information, which increases processing demands and activates system two reasoning. In the unbalanced condition, this reasoning process may be misdirected by deceptive effects of class imbalance, as AI’s overestimation of rare diseases could be interpreted as diagnostic certainty, amplifying BRN. Conversely, unbiased AI appears to improve decision outcomes by facilitating more accurate analysis within system two processing. This underscores the importance of designing AI systems that align with human cognitive tendencies, reducing the risk of compounding biases in AI-assisted decision-making.
@@
If AI biases in screening tools and symptom checkers reinforce human biases, the implications could be far-reaching. Such tools may exacerbate web-based health anxiety, also known as cyberchondria, which can cause significant distress (ryen2009cyberchondria) and in some cases lead to disorders requiring clinical attention (starcevic_cyberchondria_2019; mathes2018cyberchondria). On a societal level, cyberchondria has been associated with unnecessary doctor visits that strain healthcare systems, inflate costs, and divert resources from critical medical needs (Chambers2019Digital). Our findings highlight the potential for biased AI to amplify pre-existing human bias, reinforcing behaviours with consequences for both individual health and society.
@@
Beyond healthcare, compound biases have also been shown to negatively influence outcomes in other domains (glickmanHowHumanAI2024; winderBiasedEchoesLarge2025). This underlines the need to test different biases across contexts and to further explore their interactions (vonfeltenIsolationInteractionistPerspective2025). High-stakes areas such as hiring, legal justice, or insurance may be particularly vulnerable if compounded human-AI biases are at play. For example, in hiring, an AI system biased by gender or race disparities could interact with (un)conscious implicit bias, resulting in greater discrimination. Likewise, in legal or insurance settings, biased AI tools could mislead decision-makers and further entrench systemic inequities. These cases emphasize the importance of examining how specific AI and human bias interactions across domains.
@@
Finally, the findings of this study have implications for AI research and development. While class imbalance is a known issue for technical performance (guo2008class), our results highlight its impact on appropriate reliance, emphasizing the need for balanced datasets beyond prediction accuracy. This goal can be achieved by technical means (e.g., imbalanced learning, down-sampling) or by regulators that mandate and control class distributions and other dataset characteristics. One possible regulatory solution could be ”datasheets,” which accompany datasets and document their motivation, composition, collection process, and recommended uses (gebru2021datasheets). Coupling datasheets with explanation techniques could further support human understanding of AI predictions. However, too much information may cause cognitive overload and reduce decision quality, making it important to balance explanation density with cognitive effort (Abdul2020cogam).
@@
Overall, our study underscores the potential for biased AI to obscure reliance calibration and amplify pre-existing human biases, with consequences that extend beyond healthcare into other high-stakes domains. The HCI community must therefore actively discuss and develop design guidelines, ethical frameworks, and governance measures to prevent such outcomes. This includes designing AI systems that account for both human and AI biases jointly, striving for more accurate, reliable, and responsible AI-assisted decision-making (vonfeltenIsolationInteractionistPerspective2025).
@@ 8. Limitations and Future Research
Third, the experiment required numerous trials per participant, increasing participant burden. This design was necessary because RAIR and RSR calculations can only use half of the possible scenarios as reliance outcomes. While the extended interaction time mirrors real-world AI use, it posed practical challenges. Data cleaning addressed potential fatigue effects, and overall compliance and data quality were good. Future research could examine whether similar effects are observed with healthcare professionals (Gigerenzer1998AIDS; gigerenzer_helping_2007), particularly as AI tools become more integrated into clinical practice (Panigutti2022understanding; Levy2021assessing; Lee2021humanAIcollab).
@@ 12. Funding and Declaration of Conflicting Interests
13. Appendix
13.1. Model Selection Multinomial Logistic Regression Model
Model Equation AIC
Model 1: $\text{Disease}\sim 1$ 24013.55
Model 2: $\text{Disease}\sim\text{advice}+\text{balancing}$ 23904.88
Model 3: $\text{Disease}\sim\text{advice}\times\text{balancing}$ 23858.82
Model 4: $\text{Disease}\sim\text{advice}+\text{balancing}+\text{Trial}$ 23893.04
Model 5: $\text{Disease}\sim\text{advice}\times\text{balancing}+\text{Trial}$ 23846.97
Model 6: $\text{Disease}\sim\text{advice}\times\text{balancing}\times\text{Trial}$ 23844.73
Table 3. Model selection: Akaike Information Criterion (AIC) comparison.
Model Comparison df Deviance $p$
Model 2 vs Model 1 4 116.67 <0.001<0.001
Model 3 vs Model 2 2 50.07 <0.001<0.001
Model 4 vs Model 3 0 -34.23 n.s.
Model 5 vs Model 4 2 50.07 <0.001<0.001
Model 6 vs Model 5 6 14.25 =0.02699=0.02699
Table 4. Likelihood-Ratio test for model comparisons. If the model 4 vs model 3 comparison is left out, the final selection remained the same.
13.2. Study Procedure
Figure 6. Flowchart of the study procedure.:
13.3. Data Generation Process Illustration.
Figure 7. Illustration of the data generation process for the symptom ”runny nose” in a fictitious patient diagnosed with dengue.
1: Dengue is the disease this patient suffers from. 2: Its association with the current symptom ”runny nose” is evaluated. 3: On three out of four websites dengue is associated with a runny nose, resulting in a frequency rating of three. 4: Based on this frequency rating, a random number from the corresponding probability range (51% to 75%) is generated, yielding a value of 68% in this example. 5: This probability was used to determine symptom presence by sampling from a binomial distribution. 6: In this case, a ”one” was drawn, indicating that this particular patient with dengue experienced a runny nose. Then this process repeats and the next symptom (in this example, ”fever”) was generated for this patient.:

### v2

@@ 7.2.4. Broader Implications for AI Research and Practice
Integrating BRN within the broader framework of human biases, our findings may support the view that BRN involves deliberate processing (system two) rather than being purely automatic and intuitive (Pennycook et al., 2014; Markovits and Béghin, 2023). A plausible cognitive mechanism is that AI advice introduces additional information, which increases processing demands and activates system two reasoning. In the unbalanced condition, this reasoning process may be misdirected by deceptive effects of class imbalance, as AI’s overestimation of rare diseases could be interpreted as diagnostic certainty, amplifying BRN. Conversely, unbiased AI appears to improve decision outcomes by facilitating more accurate analysis within system two processing. This underscores the importance of designing AI systems that align with human cognitive tendencies, reducing the risk of compounding biases in AI-assisted decision-making.
@@
If AI biases in screening tools and symptom checkers reinforce human biases, the implications could be far-reaching. Such tools may exacerbate web-based health anxiety, also known as cyberchondria, which can cause significant distress (White and Horvitz, 2009) and in some cases lead to disorders requiring clinical attention (Starcevic et al., 2019; Mathes et al., 2018). On a societal level, cyberchondria has been associated with unnecessary doctor visits that strain healthcare systems, inflate costs, and divert resources from critical medical needs (Chambers et al., 2019). Our findings highlight the potential for biased AI to amplify pre-existing human bias, reinforcing behaviours with consequences for both individual health and society.
@@
Beyond healthcare, compound biases have also been shown to negatively influence outcomes in other domains (Glickman and Sharot, 2024; Winder et al., 2025). This underlines the need to test different biases across contexts and to further explore their interactions (von Felten, 2025). High-stakes areas such as hiring, legal justice, or insurance may be particularly vulnerable if compounded human-AI biases are at play. For example, in hiring, an AI system biased by gender or race disparities could interact with (un)conscious implicit bias, resulting in greater discrimination. Likewise, in legal or insurance settings, biased AI tools could mislead decision-makers and further entrench systemic inequities. These cases emphasize the importance of examining how specific AI and human bias interactions across domains.
@@
Finally, the findings of this study have implications for AI research and development. While class imbalance is a known issue for technical performance (Guo et al., 2008), our results highlight its impact on appropriate reliance, emphasizing the need for balanced datasets beyond prediction accuracy. This goal can be achieved by technical means (e.g., imbalanced learning, down-sampling) or by regulators that mandate and control class distributions and other dataset characteristics. One possible regulatory solution could be ”datasheets,” which accompany datasets and document their motivation, composition, collection process, and recommended uses (Gebru et al., 2021). Coupling datasheets with explanation techniques could further support human understanding of AI predictions. However, too much information may cause cognitive overload and reduce decision quality, making it important to balance explanation density with cognitive effort (Abdul et al., 2020).
@@
Overall, our study underscores the potential for biased AI to obscure reliance calibration and amplify pre-existing human biases, with consequences that extend beyond healthcare into other high-stakes domains. The HCI community must therefore actively discuss and develop design guidelines, ethical frameworks, and governance measures to prevent such outcomes. This includes designing AI systems that account for both human and AI biases jointly, striving for more accurate, reliable, and responsible AI-assisted decision-making (von Felten, 2025).
@@ 8. Limitations and Future Research
Third, the experiment required numerous trials per participant, increasing participant burden. This design was necessary because RAIR and RSR calculations can only use half of the possible scenarios as reliance outcomes. While the extended interaction time mirrors real-world AI use, it posed practical challenges. Data cleaning addressed potential fatigue effects, and overall compliance and data quality were good. Future research could examine whether similar effects are observed with healthcare professionals (Gigerenzer et al., 1998, 2007), particularly as AI tools become more integrated into clinical practice (Panigutti et al., 2022; Levy et al., 2021; Lee et al., 2021).